OpenStudy (anonymous):

Write the point slope form of the equation of the line passing through the point ( -12, -5) with a slope of 3.

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

The point-slope form of a linear equation looks like this: y - y1 = m*(x - x1) where m is the slope of the line and (x1,y1) is a point on the line.

Directrix (directrix):

@DeeBush96 Use that equation ^^^ to fill in for the slope and the point given in the problem. Post what you get and we will compare answers, okay?

OpenStudy (anonymous):

i think it is y + 5+ 3(x+12) @Directrix

Directrix (directrix):

Yes, I almost agree. I think there's a typo here: y + 5+ 3(x+12) --> y + 5 = 3(x + 12)

OpenStudy (anonymous):

so im rite

Directrix (directrix):

Yes, it you put an "equals" in your equation.

OpenStudy (anonymous):

Find the values of x and y. The diagram is not a scale @Directrix

Directrix (directrix):

What is true about the two angles marked in red on the attachment? It has to do with parallel lines.

Directrix (directrix):

What name is given to those two angles marked in red? The name has to do with angles formed by two parallel lines cut by a transversal.

OpenStudy (anonymous):

is it angle bisectors

Directrix (directrix):

No. They are alternate interior angles. Theorem: When two parallel lines are cut by a transversal, alternate interior angles are congruent.

OpenStudy (anonymous):

ok

Directrix (directrix):

So, both have measure 40. Now, look at the diagram attached.

OpenStudy (anonymous):

is X= 75 and Y=68

Directrix (directrix):

How did you get those?

Directrix (directrix):

Solve for y: 40 + y - 2 + 74 = 180 --> Sum of angles in triangle is 180

Directrix (directrix):

Simplify this: 40 + y - 2 + 74

OpenStudy (anonymous):

can u help me simplify this

Directrix (directrix):

40 + y - 2 + 74 = 40 - 2 + 74 + y = ? Does that help?

Directrix (directrix):

40 - 2 + 74 = ?

OpenStudy (anonymous):

112

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Okay, so 40 + y - 2 + 74 = 180 112 + y = 180 y = ?

OpenStudy (anonymous):

so do i subtract 112 from 180 @Directrix

Directrix (directrix):

Yes y = 180 - 112 y = ?

OpenStudy (anonymous):

68 what is 8

Directrix (directrix):

y = 68

OpenStudy (anonymous):

i meant what is x

Directrix (directrix):

Here is how to find x.

Directrix (directrix):

Solve for x: x -1 + 40 + 66 = 180

OpenStudy (anonymous):

is it 75

Directrix (directrix):

Guessing from the options? :)

Directrix (directrix):

Yes, 75 = x.